OpenStudy (anonymous):

Simplify: √(x^2+2x+1)

OpenStudy (saifoo.khan):

\[\sqrt{x^2+2x+1}\]Like that?

OpenStudy (anonymous):

Yes; would it be 2x^3+1 the squared?

OpenStudy (saifoo.khan):

\[\sqrt{(x+1)(x+1)}\]\[\sqrt{(x+1)^2}\]Radical cancels with square. so you are left with:\[x+1\]

OpenStudy (saifoo.khan):

No, it wouldn't be that.